OpenStudy (asianna1):

A scientist writes the equation n(h)100^e^0.25h to model the growth of a certain bacteria in a petri dish, where N represents the number of bacteria after h hours. After approximately how many hours will 450 bacteria be present? Round your answer to the nearest whole number.

OpenStudy (jack1):

hiya so you've used logarithms before, yeah?

OpenStudy (jack1):

n(h)= 100^e^0.25h 450 = 100^e^0.25h then use log and ln to solve for h